for each strong acid solution determine [H3O+] [OH-] ans pH

112. For each strong acia and pt. (a) o. Biso M HCl (6) 1.9810-14m HI (c).0.0226 m HBr (d) 1.78103 M NHINO
Answer #1

All are strong acids dissociates completely into ions . As all are monoprotic acids, the concentration of hydronium ion formed on dissociation is equal to concentration of acid. That is

[H3O+] = [acid]
pH is given as

pH = -log [H3O+]

At 25oC    [H3O+] [OH-] = 10-14

[OH-] = 10-14/[H3O+]

On the basis of above formula let us do calculation for each part

a) [H3O+] = 0.0150 M

pH = -log (0.0150) = 1.82

[OH-] = 10-14/0.0150 = 6.6 x 10-13 M

b ) Here the concentration of [H3O+] from acid is very less than the hydronium ions from water so the concentration of hydronium ion is given as [H3O+] ~ 10-6.99 M

pH = -log (10-6.99) = 6.99

[OH-] = 10-14/10-6.99= 9.78 M

c) [H3O+] = 0.0226 M

pH = -log (0.0226) = 1.64

[OH-] = 10-14/0.0226 = 4.4 x 10-13 M

a) [H3O+] = 0.0017 M

pH = -log (0.0017) = 2.77

[OH-] = 10-14/0.0017 = 5.8 x 10-12 M